# Knowledge-Guided Prompt Learning for Request Quality Assurance in Public Code Review

### v1

@@ Abstract.
Public Code Review (PCR) is an assistant to the internal code review of the development team, in the form of a public Software Question Answering (SQA) community, to help developers access high-quality and efficient review services. Current methods on PCR mainly focus on the reviewer’s perspective, including finding a capable reviewer, predicting comment quality, and recommending/generating review comments. However, it is not well studied that how to satisfy the review necessity requests posted by developers which can increase their visibility, which in turn acts as a prerequisite for better review responses. To this end, we propose a Knowledge-guided Prompt learning for Public Code Review (KP-PCR) to achieve developer-based code review request quality assurance (i.e., predicting request necessity and recommending tags subtask). Specifically, we reformulate the two subtasks via 1) text prompt tuning which converts both of them into a Masked Language Model (MLM) by constructing prompt templates using hard prompt; 2) knowledge and code prefix tuning which introduces external knowledge by soft prompt, and uses data flow diagrams to characterize code snippets. Finally, both of the request necessity prediction and tag recommendation subtasks output predicted results through an answer engineering module. In addition, we further analysis the time complexity of our KP-PCR that has lightweight prefix based the operation of introducing knowledge. Experimental results on the PCR dataset for the period 2011-2023 demonstrate that our KP-PCR outperforms baselines by 8.3%-28.8% in the request necessity prediction and by 0.1%-29.5% in the tag recommendation. The code implementation is released at https://github.com/WUT-IDEA/KP-PCR.
public code review, review necessity prediction, tag recommendation, prompt learning, external knowledge
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 8ccs: Software and its engineering Software post-development issues
@@ 1. Introduction
Code review can be conducted either within a team or by public developers. As an application form of traditional code review (Shan et al., 2022; Chen et al., 2022b; Li et al., 2023a; Rahman et al., 2017; Pandya and Tiwari, 2022), code review within a team, serves as an essential quality assurance mechanism in mature and successful commercial and Open Source Software (OSS) projects (Bosu and Carver, 2013). However, due to its labor-intensive nature, code review within the team gradually increases the time and resources spent on review activities as projects become more complex and extensive (Rigby and Bird, 2013; Rigby et al., 2014). Such cost requirements slow down the popularity of team code review (Johnson, 1998). Consequently, modern code review has evolved from a formal and rigorously regulated process (Sadowski et al., 2018; Fagan, 2011) to a less stringent practice (Baum et al., 2017a, 2016; Rigby and Bird, 2013).
Public Code Review (PCR), as an assistant to internal code review within teams, takes the form of Software Question Answering (SQA) community to assist developers in exploring high-quality and efficient review services (Hong et al., 2022). Developers have widely adopted it as a new option for code review. Research has found that the code review process not only helps to improve the overall quality of code and software systems (Ackerman et al., 1989; Baum et al., 2017b), identify defects (Baum and Schneider, 2016), but also promotes the exchange of professional and development technical knowledge (Sadowski et al., 2018). As an emerging form of service, in PCR developers publish review requests within a community for all community members to view, then the reviewers complete the review service through comments. In the PCR process implemented by the SQA community, request quality assurance for code review service mainly depends on developers. As developers are required and encouraged to submit PCR requests in the community, the necessity of these code review requests is evaluated by public via request necessity prediction (Ochodek et al., 2022; Chen et al., 2022a; Hong et al., 2022) subtask after they are posted. Meanwhile, developers always need to choose technical terms as review tags. These tags can be used to match requests with suitable reviewers via tag recommendation (Li et al., 2020; Nie et al., 2020; Xu et al., 2021; He et al., 2022; Li et al., 2023b) subtask.
Existing methods on PCR are mostly based on pre-trained language models. However, the language models lack the learning of professional knowledge vocabulary in the field of PCR. Early work regarded code snippets as non-linguistic characters that would affect the model’s effectiveness, for example, the TagDC method proposed by Li et al. (Li et al., 2020) and the PROFIT model proposed by Nie et al. (Nie et al., 2020). These works have treated code snippets as noise and directly discarded code snippets. Recent studies have discovered semantic information embedded in code snippets, but these studies mostly employ the same processing methods as they do with text (Xu et al., 2021; He et al., 2022). Overall, existing work has not fully utilized the knowledge embedded in code snippets.
To this end, we propose Knowledge-guided Prompt learning for Public Code Review (KP-PCR). Specifically, we first design a task-descriptive prompt template that restructures the request necessity prediction subtask and the tag recommendation subtask into a generative task through a knowledge-guided prompt template, and input the request body (including tags and descriptions) into the prompt construction module of a language model. We reformulate both subtasks via 1) text prompt tuning which converts two subtasks into Masked Language Model (MLM) by constructing prompt templates using hard prompt; and 2) knowledge and code prefix tuning which introduces external knowledge by soft prompt, and uses data flow diagrams to characterize code snippets. Finally, the request necessity prediction and recommendation tags output predicted results through an answer engineering module. In addition, we delve deeper into the time complexity analysis of our KP-PCR with lightweight prefix based operations to introducing knowledge, which reveals that our KP-PCR can improve task accuracy performance without compromising overall efficiency. Experimental results demonstrate that our KP-PCR outperforms baselines in terms of accuracy-related metrics across both subtasks. Furthermore, we explore the impact of prompt learning construction on the results, and use a case study to compare our KP-PCR with GPT-3.5 and GPT-4.0 considering tag recommendation.
The main contributions of this study can be summarized as follows:
  • We introduce the external knowledge from Wikipedia to enhance the representation of code snippets in PCR to address the semantic understanding of domain knowledge in developer PCR requests.
This paper extends the preliminary version of our study, as presented in (Chen et al., 2024). In comparison, this extended version (1) introduces external knowledge from Wikipedia to generate prefix vectors as soft prompt for the representation of code snippets, (2) designs the knowledge-guided prompt learning on the request necessity prediction subtask and the tag recommendation subtask to show that the knowledge guide can further improve the ability of our KP-PCR method in dealing with the code snippets, (3) comprehensively evaluates the effects of our task-descriptive prompt template to explore the influence of external knowledge on unifiedly modelling the above two subtasks, (4) analyzes the time complexity to prove that our KP-PCR can improve the accuracy of request necessity prediction and tag recommendation while maintaining the overall efficiency of our KP-PCR method, and (5) conducts a case study to investigate the performance of our prompt template on large language models. Additionally, we have examined the challenges of how to design hard and soft prompt templates and delved further into the originally explored unified modelling in the initial study. More details on the extension points compared to our preliminary study in (Chen et al., 2024) can be found in Section 7.5.
The remainder of this paper is structured as follows. Section 2 presents a motivating example, and Section 3 introduces the task addressed in this paper. Section 4 details the proposed method. Section 5 provides the experimental setup, and Section 6 discusses the experiment results and presents a case study to explore the utilization of external knowledge by our KP-PCR. Section 7 describes the related work. Section 8 concludes this work with future directions.
2. Motivating Example
Public Code Review (PCR) is developed in the Software Question Answering (SQA) community. Researchers are increasingly exploring ways to achieve high-quality and efficient review services (Hong et al., 2022; Ackerman et al., 1989; Baum et al., 2017b; Baum and Schneider, 2016; Sadowski et al., 2018). A review request consists of several parts, as shown in Fig. 1, including title, request description (including text and code snippets), quality, and tags.
@@ 2. Motivating Example
Since PCR differs from team internal code review in service form, the quality of the PCR process mainly depends on the developers who write and submit the review requests rather than the reviewers (Xu et al., 2021; Li et al., 2023b). Existing research is highly effective for review quality, but there is still a room for improvement. Due to the lack of pre-learning of domain knowledge by pre-trained language models, existing methods inadequately learn the domain knowledge implicit in request description.
One major characteristic of the request body is the abundance of code snippets, which is particularly significant for PCR. Specifically, for the task of assessing the necessity in code reviews, it is required by the rules of PCR websites that the code snippets in the requests should constitute runnable and correct code; otherwise, the code snippets fail to meet the review standards. As for the task of tag recommendation, even large language models that currently excel in various aspects still showed pay attention on professional knowledge tags. These factors affecting task performance are closely related to the professional knowledge contained in both text and code snippet semantics.
Figure 2. An Example of the Subject Matter Involving Professional Knowledge in a Request.
[…]
The code snippet can be clearly identified to the java language information.
Code snippets are beneficial for understanding and distinguishing the requests of public code reviews. Taking a specific review request as an example, as shown in Fig. 2, code snippets can clearly identify Java language information. By selecting processing and representation methods that are consistent with the semantic structure of the code snippets, the variable flow expression of the code snippets can be better understood, revealing the information embedded in the logic of the Java programming language. However, the professional knowledge and logical expressions contained in these code snippets may not fully pre-learned by text-based language models.
The text of review requests may contain some vocabulary that is not present in the language model. When tokenizing the text, the language model may split these out-of-vocabulary words into subwords or letters, which limit the model’s performance in completing public code review tasks. Table 1 illustrates an example of applying specialized knowledge in the context of the tag recommendation subtask. As shown in this example, a review request contains the abbreviation “OOP” to describe the content, which is not covered in the language model. However, its meaning is semantically consistent with the label “Object oriented”. Since the language model has not learned “OOP”, it may struggle to associate the abbreviation “OOP” with the label “Object oriented”. In contrast, in knowledge repositories such as Wikipedia, the definition content would link these two terms. Thus, this has motivated us to carry out this study that aims to bridge this gap.
@@ 3.1. Public Code Review Process
First, the public code review process will be introduced through an example of public code review. As shown in Fig. 3, a review process can be abstracted into 5 steps, labeled as ①-⑤.
@@
Code review typically involves two roles: the developer who submits the code request and the reviewer who provides suggestions and comments. In Step ①, the developer writes and submits a review request. In Step ②, practitioners in the community review and evaluate the requests to allocate necessary feedback. Subsequently, in Step ③, developers are asked to select several relevant technical terms to accompany their request. These tags are then used in Step ④ to match the request with appropriate reviewers, who provide their feedback on the code in Step ⑤.
In a review request, there are often many changes and optimizations needed in different parts of the code. Reviewers need to spend a significant amount of time reviewing all the changes. However, it is often the case that most changes are minor and do not require comments or suggestions. Li et al. (Li et al., 2022) initially considered the task of assessing code change quality in the internal team code review process, which involves predicting whether a code change is of high quality and ready to be accepted in the review process. This work enhances the efficiency and quality of the review process by pre-evaluating before the request is reviewed. Therefore transitioning to the public code review process, there is a task need for predicting request necessity. Request necessity prediction learns from the existing conventions of public code review and community feedback to determine the standards for submitting review requests, thereby ensuring review quality and reducing the time cost of unnecessary requests.
@@ 4.1. Framework Overview
The knowledge-guided prompt learning module consists of text prompt, code prompt and knowledge-guided prompt. First, the request text is input into prompt learning module. The text prompt reconstructs the request text into a generative task, and the code prompt refactors the code snippets in text into the form of data flow graph. These processes are described in more detail in Section 4.2. The request is then matched with the external knowledge base from Wikipedia to get the knowledge guidance, which is then fed into the model in the form of prefix vectors. Further, the obtained prompt template is input into the Masked Language Model (MLM) to get the predicted word list. Then the answer engineering module maps the predicted words to the label list to get the final labels.
Figure 4. The Framework of KP-PCR
@@ 4.2.1. Text Prompt
Finally, with a guiding prompt “Request: ”, the model is directed to explicitly recognize the request content, and concatenate the title $title$, description $des$ (including text content $text$ and code snippet content $code$), and the template $T\left(\cdot\right)$ together. After concatenation, we obtain a text input $req_{i}$ for the model, as shown in Equation 1:
@@ 4.2.2. Code Prompt
This section mainly introduces the processing of refactoring code snippets into data flow graph based vector. That is used as part of soft prompt.
Code snippets inherently possess a graph structure that provides valuable semantic information for code comprehension. To better capture this information, Guo et al. (Guo et al., 2021) proposed using data flow graphs to analyze code and represent its semantic information. In previous research on code representation, the parsing of code typically involved using Abstract Syntax Trees (ASTs). In contrast, this graph-based data flow structure effectively avoids the unnecessary depth hierarchy that ASTs may introduce.
Figure 5. A Code Snippet Processing Example
@@
the language of the code snippet is identified.the abstract syntax tree is further parsed into the form of a data flow graph
As shown in Fig. 5, the processing of refactoring code snippets into a data flow graph based vector involves two steps. Firstly, the language of the code snippet is identified, as each programming language has different coding styles and syntax. Therefore, it is necessary to parse the code snippet for each language, extract important nodes and variable information in the form of an abstract syntax tree (AST) after obtaining the language category. Secondly, the abstract syntax tree is further parsed into the form of a data flow graph where nodes are used to record the variable nodes of the code snippet, and edges represent the flow of information between nodes. For example, assigning the value of $b$ to $a$ would be parsed into two nodes, $a$ and $b$, with an edge containing the value of $b$ representing the flow of data.

### v3

@@ Abstract.
Public Code Review (PCR) is developed in the Software Question Answering (SQA) community, assisting developers in exploring high-quality and efficient review services. Current methods on PCR mainly focus on the reviewer’s perspective, including finding a capable reviewer, predicting comment quality, and recommending/generating review comments. However, it is not well studied that how to satisfy the review necessity requests posted by developers which can increase their visibility, which in turn acts as a prerequisite for better review responses. To this end, we propose Knowledge-guided Prompt learning for Public Code Review (KP-PCR) to achieve developer-based code review request quality assurance (i.e., predicting request necessity and recommending tags subtask). Specifically, we reformulate the two subtasks via 1) text prompt tuning which converts both of them into a Masked Language Model (MLM) by constructing prompt templates using hard prompt; and 2) knowledge and code prefix tuning which introduces knowledge guidance from fine-tuned large language models by soft prompt, and uses program dependence graph to characterize code snippets. Finally, both of the request necessity prediction and tag recommendation subtasks output predicted results through an answer engineering module. In addition, we further analysis the time complexity of our KP-PCR that has lightweight prefix based the operation of introducing knowledge guidance. Experimental results on the PCR dataset for the period 2011-2023 demonstrate that our KP-PCR outperforms baselines by 2.3%-8.4% in the request necessity prediction and by 1.4%-6.9% in the tag recommendation. The code implementation is released at https://github.com/WUT-IDEA/KP-PCR.
public code review, review necessity prediction, tag recommendation, prompt learning, knowledge enhancement
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 8ccs: Software and its engineering Software post-development issues
@@ 1. Introduction
Code review can be conducted either within a team or by public developers. As an application form of traditional code review (Shan et al., 2022; Chen et al., 2022b; Li et al., 2023a; Rahman et al., 2017; Pandya and Tiwari, 2022), code review within a team, serves as an essential quality assurance mechanism in mature and successful commercial and Open Source Software (OSS) projects (Bosu and Carver, 2013). However, due to its labor-intensive nature, code review within team gradually increases the time and resources spent on review activities as projects become more complex and extensive (Rigby and Bird, 2013; Rigby et al., 2014). Therefore, beside the popularity of team code review (Johnson, 1998), modern code review has evolved from a formal and rigorously regulated process (Sadowski et al., 2018; Fagan, 2011) to a less stringent practice (Baum et al., 2017a, 2016; Rigby and Bird, 2013).
Public Code Review (PCR) is a type of modern code review task developed in the Software Question Answering (SQA) community, which helps developers explore efficient and high-quality review services. Unlike traditional code review, PCR is typically directed towards a broader developer community, where users can submit review requests without being tied to a specific project context. This openness enables PCR to gather diverse perspectives and expertise on code, but the lack of project structure often leads to submissions that lack context or clear background information, introducing challenges in management and quality control of code review. Existing research on code review has already recognized the importance of quality assurance mechanisms (Baum et al., 2017b; Baum and Schneider, 2016; Sadowski et al., 2018). Therefore, this issue should also be given attention in public code review tasks. In open environments lacking project structure, appropriate quality control not only helps guide users to submit clearer and more standardized review requests, but also provides safeguards for identifying potential issues and facilitating efficient collaboration.
In the PCR process implemented by the SQA community, request quality assurance for code review service mainly depends on developers. Developers are required and encouraged to submit PCR requests in the community. A review request [^1] https://codereview.stackexchange.com/questions/284179/proper-implementation-of-signal-handler-and-multithreading-pthread consists of several parts, as shown in Fig. 1, including title, request description (text and code snippets), quality, and tags. The necessity of these code review requests is evaluated by public via request necessity prediction (Chen et al., 2024) subtask after the requests are posted. Meanwhile, developers always need to choose technical terms as review tags. These tags can be used to match requests with suitable reviewers via tag recommendation (Li et al., 2023b; Zhou et al., 2019) subtask. Since the quality and responsiveness of public code reviews typically rely on manual efforts by developers, which can be labor-intensive, the demand for transitioning from manual reviews to intelligent services has gradually emerged.
@@
Existing methods on PCR are mostly based on pre-trained language models. However, these models typically exhibit generalized capabilities, and their understanding of domain-specific knowledge in the PCR field, such as code-related semantic patterns, still has room for improvement. Early work regarded code snippets as non-linguistic characters, which impeded the model’s performance in understanding technical content. For example, the TagDC method proposed by Li et al. (Li et al., 2020) and the PROFIT model proposed by Nie et al. (Nie et al., 2020) treated code snippets as irrelevant information and discarded them altogether. Recent studies have discovered valuable semantic information embedded in code snippets, but these studies mostly employed the same processing methods as they do with text (Xu et al., 2021; He et al., 2022). Knowledge enhancement can effectively address the gaps in existing models when handling programming terms and error patterns. For instance, entries on algorithms, programming languages, and common programming errors can help the model better understand the functionality and potential issues of complex code snippets, thus improving its performance in PCR tasks.
To this end, we propose Knowledge-guided Prompt learning for Public Code Review (KP-PCR). Specifically, we first design a task-descriptive prompt template that restructures the request necessity prediction subtask and the tag recommendation subtask into a generative task through a knowledge-guided prompt template, and input the request body (including tags and descriptions) into the prompt construction module of a language model. We reformulate both subtasks via 1) text prompt tuning which converts two subtasks into Masked Language Model (MLM) by constructing prompt templates using hard prompt; and 2) knowledge and code prefix tuning which introduces knowledge guidance from fine-tuned DeepSeek by soft prompt, and uses program dependence graph to characterize code snippets. Finally, the request necessity prediction and recommendation tags output predicted results through an answer engineering module. In addition, we delve deeper into the time complexity analysis of our KP-PCR with lightweight prefix based operations to introducing knowledge, which reveals that our KP-PCR can improve task accuracy performance without compromising overall efficiency. Experimental results demonstrate that our KP-PCR outperforms baselines in terms of accuracy-related metrics across both subtasks. Furthermore, we explore the impact of prompt learning construction and knowledge guidance sources on the results, and use a case study to compare our KP-PCR with Llama3, CodeLlama and DeepSeek.
The main contributions of this study can be summarized as follows:
  • We introduce knowledge guidance from fine-tuned DeepSeek to enhance the model’s semantic understanding of domain knowledge in developers’ PCR requests.
This paper extends the preliminary version of our study, as presented in (Chen et al., 2024). In comparison, this extended version (1) introduces knowledge guidance from fine-tuned DeepSeek to generate prefix vectors as soft prompt for the representation of code snippets, (2) changes the code parsing mode to PDG to get more control dependency information, (3) analyzes the time complexity to prove that our KP-PCR can improve the accuracy of request necessity prediction and tag recommendation while maintaining the overall efficiency of our KP-PCR method, (4) comprehensively evaluates the effects of our task-descriptive prompt template to explore the influence of knowledge guidance and code parsing on unifiedly modelling the above two subtasks, and (5) conducts a case study to compare our KP-PCR with Llama3, CodeLlama and DeepSeek. Additionally, we have examined the challenges of how to design hard and soft prompt templates and delved further into the originally explored unified modelling in the initial study. More details on the extension points compared to our preliminary study in (Chen et al., 2024) can be found in Section 7.6.
The remainder of this paper is structured as follows. Section 2 presents a motivating example, and Section 3 introduces the task addressed in this paper. Section 4 details the proposed method. Section 5 provides the experimental setup, and Section 6 discusses the experiment results and presents a case study to explore the utilization of knowledge guidance by our KP-PCR. Section 7 describes the related work. Section 8 concludes this work with future directions.
2. Motivating Example
Public Code Review (PCR) differs from internal team code reviews in its service format, as the quality of the PCR process primarily depends on the developers who write and submit review requests, rather than on the reviewers themselves. While existing research has been effective in improving review quality, there remains significant room for improvement. One of the limitations of the current approach is that pre-trained language models still require improvement in their ability to learn domain-specific knowledge. The text of review requests may also include vocabulary that is absent from the language model’s training data. During tokenization, the model may split these out-of-vocabulary words into subwords or individual characters, which limits its ability to accurately process the text and complete PCR tasks.
Table 1 provides an example from the tag recommendation subtask, in which the review request includes the abbreviation “OOP” to describe a concept. However, this term is not covered by the language model, and the model may struggle to connect “OOP” with its full form, “Object Oriented Programming”. In contrast, knowledge guidance provides definitions that link these terms, helping to resolve such discrepancies.
To address this limitation, we propose fine-tuning a large pre-trained model, such as DeepSeek. By fine-tuning DeepSeek-7B on Wikipedia’s extensive collection of programming terminology, technical definitions, and domain-specific concepts, we can improve the model’s ability to understand and generate code-related content more accurately. This approach allows the model to internalize valuable knowledge during the training process, ensuring that the learned domain-specific knowledge is embedded directly within the model’s parameters.
@@ 3.1. Public Code Review Process
First, the public code review process will be introduced through an example of public code review. As shown in Fig. 2, a review process can be abstracted into 5 steps, labeled as ①-⑤. Code review typically involves two roles: the developer who submits the code request and the reviewer who provides suggestions and comments. In Step ①, the developer writes and submits a review request. In Step ②, practitioners in the community review and evaluate the requests to allocate necessary feedback. Subsequently, in Step ③, developers are asked to select several relevant technical terms to accompany their request. These tags are then used in Step ④ to match the request with appropriate reviewers, who provide their feedback on the code in Step ⑤.
@@
In a review request, there are often many changes and optimizations needed in different parts of the code. Reviewers need to spend a significant amount of time reviewing all the changes. However, it is often the case that most changes are minor and do not require comments or suggestions. Li et al. (Li et al., 2022) initially considered the task of assessing code change quality in the internal team code review process, which involves predicting whether a code change is of high quality and ready to be accepted in the review process. This work enhances the efficiency and quality of the review process by pre-evaluating before the request is reviewed. Therefore, transitioning to the public code review process, there is a task need for predicting request necessity. Request necessity prediction learns from the existing conventions of public code review and community feedback to determine the standards for submitting review requests, thereby ensuring review quality and reducing the time cost of unnecessary requests.
@@ 4.1. Framework Overview
The knowledge-guided prompt learning module consists of text prompt, code prompt and knowledge-guided prompt. First, the request text is input into prompt learning module. The text prompt reconstructs the request text into a generative task, and the code prompt refactors the code snippets in text into the form of program dependence graph. These processes are described in more detail in Section 4.2. The request is then fed into the fine-tuned DeepSeek. The text generated by DeepSeek is used as knowledge guidance, which is then fed into the model in the form of prefix vectors. Further, the obtained prompt template is input into the Masked Language Model (MLM) to get the predicted word list. Then the answer engineering module maps the predicted words to the label list to get the final labels.
Figure 3. The Framework of KP-PCR.
@@ 4.2.1. Text Prompt
Finally, with a guiding prompt “Request: ”, the model is directed to explicitly recognize the request content, and concatenate the title titletitle, description desdes (including text content texttext and code snippet content codecode), and the template $T\left(\cdot\right)$ together. After concatenation, we obtain a text input $req_{i}$ for the model, as shown in Equation 1:
@@ 4.2.2. Code Prompt
This subsection systematically elaborates on the methodology for constructing prefix vectors based on program dependence analysis.
In the label recommendation task, labels are often closely related to the behavior of code under different control flow paths. Similarly, in the necessity prediction task, it is often essential to understand both the control logic and the data dependencies within the surrounding context. Program Dependence Graph (PDG) captures both control dependencies and data dependencies, allowing for a comprehensive representation of the program execution logic. By hierarchically parsing code Data Flow Graph (DFG) and Control Flow Graph (CFG), as shown in Figure 4, we ultimately fuse these structures to build a Program Dependence Graph (PDG), enabling formal semantic representation of code.
Firstly, the language of the code snippet is identified, as each programming language has different coding styles and syntax. Therefore, it is necessary to parse the code snippet for each language, extract important nodes and variable information in the form of an abstract syntax tree (AST) after obtaining the language category. Secondly, the abstract syntax tree is further parsed into the form of DFG and CFG. When converting incomplete code snippets, technique such as filling in gaps with placeholder nodes is used to handle incomplete connections.
For a code snippet $C$, its DFG is defined as a directed graph $G_{dfg}=(V_{dfg},E_{dfg})$. Node set $V_{dfg}=\{v|v\in\text{Variables}(C)\cup\text{Operations}(C)\}$ represents variables and operators. Edge set $E_{dfg}=\{\left(v_{i},v_{j}\right)|v_{j}\rightarrow_{dfg}v_{i}\}$, where $v_{j}\rightarrow v_{i}$ denotes a data dependence from $v_{i}$ to $v_{j}$.
The CFG of code snippet $C$ is defined as a directed graph $G_{cfg}=(V_{cfg},E_{cfg})$. Node set $V_{cfg}=\{b|b\in\text{BasicBlocks}(C)\}$, where basic blocks are maximal sequences of branch-free instructions. Edge set $E_{cfg}=\{(b_{i},b_{j})|b_{i}\rightarrow_{cfg}b_{j}\}$, where $b_{i}\rightarrow_{cfg}b_{j}$ denotes immediate control flow from $b_{i}$ to $b_{j}$.
Finally, duplicate nodes between the DFG and CFG are merged to generate PDG with comprehensive data and control dependency representation.
Figure 4. A Code Snippet Processing Example.
@@
the language of the code snippet is identified.the abstract syntax tree is further parsed into the form of a PDG